Consider the following numerical example of the simple keynesian model with no government spending ,taxes or foreign sector (all figures in R millions):
C=100 + 0,9Y
I=50
Answer the following questions.
1-calculate the equilibrium level of output.
2-in equilibrium , what is the value of consumption spending ?use this number to verify that the sum of C and I in equilbrium equals the value for equilibrium output you obtained above.
3-suppose the level of output that creates full employment in the economy is 1800.using the multipler ,determine the level of investment spending that would create full employment in this economy.
1
Expert's answer
2019-08-13T08:56:20-0400

1. In equilibrium Y = C + I = 100 + 0.9Y + 50,

0.1Y = 150,

Y = 1500.

2. In equilibrium the value of consumption spending is C = 100 + 0.9×1500 = 1450.

3. If the level of output that creates full employment in the economy is 1800, then we need increase in GDP of 300 to close the recessional gap. The multipler is 1/(1 - 0.9) = 10, so the level of investment spending that would create full employment in this economy is 300/10 = 30.